# Does DetectGPT Fully Utilize Perturbation? Bridge Selective Perturbation to Fine-tuned Contrastive Learning Detector would be Better

### v3

@@ Abstract
The burgeoning generative capabilities of large language models (LLMs) have raised growing concerns about abuse, demanding automatic machine-generated text detectors. DetectGPT Mitchell et al. (2023), a zero-shot metric-based detector, first introduces perturbation and shows great performance improvement. However, in DetectGPT, random perturbation strategy could introduce noise, and logit regression depends on threshold, harming the generalizability and applicability of individual or small-batch inputs. Hence, we propose a novel fine-tuned detector, Pecola, bridging metric-based and fine-tuned detectors by contrastive learning on selective perturbation. Selective strategy retains important tokens during perturbation and weights for multi-pair contrastive learning. The experiments show that Pecola outperforms the state-of-the-art by 1.20% in accuracy on average on four public datasets. And we further analyze the effectiveness, robustness, and generalization of the method.
@@ 1 Introduction
Machine-generated text (MGT) detection is to discriminate MGT from human-written texts (HWT), preventing abuse of large language models (LLMs), including academic misconduct Vasilatos et al. (2023), spam synthesis Dou et al. (2020), untrustworthy news Zellers et al. (2019), etc. Currently, existing MGT detection methods can be mainly classified into two genres Wu et al. (2023), i.e., fine-tuned methods Liu et al. (2023); Hu et al. (2023); Verma et al. (2023); OpenAI (2023) and zero-shot metric-based methods Gehrmann et al. (2019); Mitchell et al. (2023); Yang et al. (2023); Bao et al. (2024). In general terms, fine-tuned detector methods can achieve better accuracy than zero-shot metric-based methods, especially generalizable to black-box generators, but are more costly during data collection, fine-tuning, and running. On the other hand, zero-shot metric-based methods show better interpretability than fine-tuned ones.
@@
However, DetectGPT still has three significant defects: (i) DetectGPT’s reliance on the logit regression module’s threshold compromises its generalization in zero-shot settings and limited to large batch input, fails on individual inputs. (ii) DetectGPT does not fully utilize the perturbation. As a metrics-based method, it only considers the probability difference caused by perturbation, which is overly simplified and slightly indistinguishable. Perturbation should indeed be a stronger augment that carries implicit language pattern information. (iii) DetectGPT perturbs the original sample randomly and unrestricted, which could introduce more noise and negatively impact the performance Kim et al. (2022). For example, Liu et al. (2023) find entity-relationship plays a role in the detection, which might be destroyed in random perturbation of DetectGPT.
In this paper, we thus propose a Perturbation-based Contrastive Learning model, Pecola, for MGT detection, toward the defects with two stages, i.e., Selective Strategy Perturbation (Sec. 3.1) and Token-Level Weighted Multi-Pairwise Contrastive Learning (Sec. 3.2). Firstly, Selective Strategy Perturbation is a token-level rewriting method with restrictions on modifying important texts Campos et al. (2020) to reduce noise. The motivation is to simulate the human behavior of modification Verma and Lee (2017); Fetaya et al. (2020); Wang et al. (2019). The perturbation strategy consists of tokens removing and substitution, as shown in Fig. 1. The experiments show that our Selective Strategy Perturbation method can improve the performance of both metrics-based (i.e., DetectGPT) and model-based methods. Secondly, we use a Multi-Pairwise Contrastive Learning model to process the perturbed texts. Different from the logit regression module in DetectGPT, the trained model is generalizable without any threshold setting, and it can deal with individual inputs. Moreover, by utilizing multi-pairwise contrastive learning, the model could better utilize perturbation to focus on the language pattern gap between HWT and MGT. The importance weight from the perturbation stage is also reused as contrastive learning weight. Notably, by using contrastive learning, Pecola is a strong few-shot fine-tuning method, which effectively bridges and integrates metric-based and fine-tuned detector genres. Finally, extensive experiments show Pecola is significantly superior to baseline and SOTA methods on four datasets, Pecola improves by 1.20% to SOTA on average under few-shot settings, surpassing the latest methods by 3.84% among metric-based detectors and by 1.62% among fine-tuned detectors. Further experiments show that Pecola is as well better at generalization, robustness, and effectiveness.
@@
Outperformance: Our detector Pecola outperforms all eight compared models on four public datasets. And Pecola is more robust to the choice of base model and filling model. Furthermore, we prove its generalization ability across domains and genres of data.
@@ 2 Related Work
Machine-generated Text Detection. While fine-tuned detectors have proven effective for MGT detection Wahle et al. (2022); Hu et al. (2023), the requirement for annotated datasets poses a significant challenge due to the proliferation of unchecked, high-quality generated texts. To address this challenge, DetectGPT Mitchell et al. (2023) and Fast-DetectGPT Bao et al. (2024) have demonstrated strong performance in white-box zero-shot settings. Similarly, CoCo Liu et al. (2023) is designed to detect MGT with low resource annotations, utilizing a coherence-based contrastive learning model. Recently, the Watermark method Kirchenbauer et al. (2023) has been introduced to mitigate the risk associated with unchecked MGTs by embedding imperceptible signals within text outputs. In contrast to previous methods, our approach integrates data perturbation with contrastive learning, with a particular emphasis on reducing reliance on mask-filling models and enhancing performance in few-shot scenarios.
@@
Perturbation. Data perturbation methods find frequent application in text classification tasks Gao et al. (2022); Shum et al. (2023), which is commonly employed through the technique of consistency regularization Xie et al. (2020); Chen et al. (2020). Nevertheless, in MGT detection, previous perturbation methods have exhibited certain limitations. For instance, they often resort to randomly selecting target tokens for synonym replacement Wang et al. (2018), deletion, or insertion Wei and Zou (2019), and fine-tuning pre-trained models to fill text spans of variable lengths Gao et al. (2022). While these methods do enhance text diversity, the indiscriminate replacement of tokens without guided rules can lead to the generation of less reliable texts. These limitations motivate us to devise data perturbation methods tailored for MGT detection. Our approach, guided selectively, aims to better represent meaningful recombination spaces while preserving the inherent semantic features of the text, ultimately enhancing the diversity of samples.
@@
Constrastive Learning. Contrastive learning is an effective solution method to the issues of relying solely on cross-entropy classification loss leading to a lack of robustness and suboptimal generalization Tack et al. (2020); Hu et al. (2023). In limited labeled data task Gunel et al. (2021), introduce a robust contrastive learning method to capture the similarities between the same instances in the representation space, while separating those of different classes. Similarly, out-of-distribution (OOD) usually leads to severe semantic shift issues during inference, prompting another approach based on margin contrastive learning Zhou et al. (2021). Differently, our method focuses more on the changes of the rephrase space in data distribution after perturbation, and strives to reduce reliance on the mask-filling models in few-shot learning.
@@ 3 Methodology
As shown in Fig. 2, the workflow of Pecola mainly consists of two stages: Selective Strategy Perturbation and Supervised Contrastive Learning for fine-tuning PLM, which enjoy the advantage of metric-based and model-based detectors, respectively.
@@ 3.1 Selective Strategy Perturbation
Token Importance Assessment. To accurately assess the significance of tokens within the text and mitigate information loss stemming from random masking, we expand upon the YAKE algorithm Campos et al. (2020) to operate at the token level. The YAKE algorithm builds upon certain assumptions Machado et al. (2009), which posit that the importance of a candidate word decreases as the richness of the vocabulary surrounding it on both sides increases. This fundamental assumption remains applicable when processing text at the token level, i.e., token importance assessment.
@@
Mask Position Selection. After getting the important tokens set of each text input , we use special token [MASK] to replace some of the tokens in the text input to construct masked perturbation input . In order to relieve the information loss caused by masking perturbation, we add regularization to the traditional random masking method and use a selective masking strategy to prevent important tokens from being masked.
@@
where is the mask ratio, and represents an indicator function with a value of if and only if the condition is satisfied, otherwise, it is . Then we gather all masked perturbation inputs and include them in the training set to give the model masked perturbation improving model robustness.
Mask-Filling. Additionally, we utilize the current PLMs, e.g., T5 Raffel et al. (2020) or RoBERTa Liu et al. (2019) etc., to fill the masked perturbation inputs and create the filled perturbation inputs . Similar to before, we include all filled perturbation inputs in the training set and obtain the final training set .
@@ 3.2 Token-Level Weighted Multi-Pairwise Contrastive Learning
This adaptive margin ensures that despite data Perturbation during training, the model is steered to maintain discriminative embeddings. Then we get the following contrastive loss as:
@@
| Dataset | Metric | Shot | RoBERTa | GLTR | CE+SCL | CE+Margin | IT:Clust | CoCo | DetectGPT | Fast-Detect. | Pecola |
@@ 4.2 Comparison Models
RoBERTa Liu et al. (2019), supervised methods via standard fine-tuning pre-trained language models as classifiers. We use RoBERTa-base (135M). GLTR Gehrmann et al. (2019), a metric-based detector and based on next-token probability. We follow the setting of Guo et al. (2023), utilizing the Test-2 feature. For a fair comparison with fine-tuning methods, we first use the few-shot training samples to settle the threshold and adapt the fixed threshold in the test set.111The base model of GLTR is chosen based on the generator of the dataset: for GPT-2 and Grover datasets, we use GPT-2 Small (124M); and for GPT-3.5 and HC3 datasets, we use GPT-J (6B) Wang (2021), which is the best open-source model to simulate ChatGPT and GPT-3.5 empirically.
@@
IT:Clust Shnarch et al. (2022), a general text classification method that employs unsupervised clustering as an intermediate for fine-tuning pre-trained models, utilizing RoBERTa-base (110M).
@@
DetectGPT Mitchell et al. (2023), a zero-shot metric-based MGT detector, using T5-large Raffel et al. (2020) to perturb texts. Same as GLTR, we fix the threshold.222For all four datasets (including HC3 and GPT-3.5 datasets), we use GPT-2 Small (124M) as the base model to calculate the likelihood. The reason is Mireshghallah et al. (2023) find that small model is better black-box detector for DetectGPT.
@@ 4.3 Performance Comparison
Moreover, the unsupervised learning methods tend to show better performance in extremely few shot scenarios. Unsurprisingly, unsupervised methods do not see a notable performance improvement with the increase in the number of training samples, which causes them to outperform on the fewest shot settings initially but soon be surpassed. As for the deception of generators, Grover appears to be the hardest to detect while other models are relatively "honest" to detectors. It might have originated from the adversarial training strategy of Grover, while the bulit-in detector module adversarially shifts the LLM’s detectable features. More interestingly, advanced language models show a weaker ability to cheat detectors. Most detectors achieve around 98% in accuracy on the GPT-3.5 and HC3 datasets, which is consistent with the conclusion from Liu et al. (2023); Chen et al. (2023). We hypothesize that the easy-to-detect nature may originate from the lack of semantics diversity in GPT-3.5 and ChatGPT as they are using RLHF Kirk et al. (2023).
@@ 4.4 Ablation Study
To illustrate the effectiveness of the Pecola components, we study the ablation experiments on the Selective Strategy Perturbation stage and the Contrastive Learning stage on the 64-example GPT-2 dataset. We also demonstrate the Scalability of Pecola in Appendix C.
| Method | Acc | F1 |
| w/o. mask | 78.00 | 77.93 |
| w/o. mask-fill | 77.78 | 77.72 |
| w/o. mask.CL | 75.80 | 75.23 |
| w/o. mask-fill. CL | 75.56 | 75.10 |
| w/o. CL | 76.60 | 76.22 |
| w/o. | 78.02 | 77.93 |
@@
Ablation on Contrastive Learning. It primarily investigates the impact of CE and contrastive loss. (i) w/o. CL refers to the model larking weighted contrastive learning; (ii) w/o. refers to the model including contrastive learning but larking weight.
@@
| Model | RoBERTa | Pecola | ||
| Metric | Acc | F1 | Acc | F1 |
| Original | 74.41 | 73.91 | 78.92 | 78.88 |
| Delete | 71.77(-2.640) | 70.42(-3.490) | 77.28(-1.640) | 77.06(-1.820) |
| Repeat | 64.69(-9.720) | 61.74(-12.17) | 69.74(-9.180) | 67.87(-11.01) |
| Insert | 50.75(-23.66) | 36.44(-37.47) | 57.61(-21.31) | 49.29(-29.59) |
| Replace | 52.04(-22.37) | 39.48(-34.43) | 57.25(-21.67) | 48.89(-29.99) |
| Average | 59.81 (-14.60) | 52.02 (-21.89) | 65.47 (-13.45) | 60.78 (-18.10) |
@@ 4.5.1 Model Qualities
Analysis on Robustness. To validate the robustness of Pecola in the few-shot learning settings, we apply four post hoc perturbation operations for each token in the test dataset randomly, i.e., deletion, replacement, insertion, and repetition. As indicated in Table 3, for each perturbation method employed, our decline rate is consistently lower compared to the baseline RoBERTa. On average, Pecola maintains a 5.66% higher accuracy and an 8.77% superior F1-score. Specifically, in the deletion method, where we introduce a 15% random perturbation, it is noteworthy that the accuracy of Pecola utilizing the Selective Strategy Perturbation method experiences a mere 1.64% decrease, underscoring its remarkable robustness.
Analysis on Affinity. Affinity pertains to alterations in data distribution resulting from perturbations, quantified by observing the fluctuations in accuracy. We demonstrate the superiority of the selective masking method over the random masking method using the Affinity metric, following the setting of DetectGPT. We applied a 15% mask proportion with a span of 2 tokens on the test dataset and simultaneously employed T5-Large Raffel et al. (2020) as the mask-filling model. We trained RoBERTa-base and Pecola on the 64-example GPT2 dataset. As shown in Table 4, in comparison to the random masking perturbation method utilized in DetectGPT, we observe a 1.92% and 0.49% increase in Affinity when employing the selective masking method. Additionally, the mask-filling method yields affinity improvements of 3.38% and 1.32% for RoBERTa and Pecola models, respectively. These results illustrate that the Selective Multi-Strategy Perturbation method introduced in this paper effectively preserves more distinguishable features between MGTs and HWTs.
| Model | RoBERTa | Pecola |
| Random Mask | -2.64 | -1.64 |
| Selective Mask | -0.72 | -1.15 |
| Mask-Filling | -4.72 | -2.66 |

### v4

# Does DetectGPT Fully Utilize Perturbation? Bridging Selective Perturbation to Fine-tuned Contrastive Learning Detector would be Better

@@ Abstract
The burgeoning generative capabilities of large language models (LLMs) have raised growing concerns about abuse, demanding automatic machine-generated text detectors. DetectGPT Mitchell et al. (2023), a zero-shot metric-based detector, first introduces perturbation and shows great performance improvement. However, in DetectGPT, the random perturbation strategy could introduce noise, and logit regression depends on the threshold, harming the generalizability and applicability of individual or small-batch inputs. Hence, we propose a novel fine-tuned detector, Pecola, bridging metric-based and fine-tuned methods by contrastive learning on selective perturbation. Selective strategy retains important tokens during perturbation and weights for multi-pair contrastive learning. The experiments show that Pecola outperforms the state-of-the-art (SOTA) by 1.20% in accuracy on average on four public datasets. And we further analyze the effectiveness, robustness, and generalization of the method. 111The code and datasets are released at https://github.com/lsc-1/Pecola.
@@ 1 Introduction
Machine-generated text (MGT) detection is to discriminate MGT from human-written texts (HWT), preventing abuse of large language models (LLMs), including academic misconduct Vasilatos et al. (2023), spam synthesis Dou et al. (2020), untrustworthy news Zellers et al. (2019), etc. Currently, existing MGT detection methods can be mainly classified into two categories Wu et al. (2023a); Wang et al. (2024), i.e., fine-tuned methods Liu et al. (2023); Hu et al. (2023); Verma et al. (2023); OpenAI (2023); Mao et al. (2024) and zero-shot metric-based methods Gehrmann et al. (2019); Mitchell et al. (2023); Yang et al. (2023); Bao et al. (2024); Wu et al. (2023b). In general terms, fine-tuned detector methods can achieve better accuracy than zero-shot metric-based methods, especially generalizable to black-box generators, but are more costly during data collection, fine-tuning, and running, in most cases. On the other hand, zero-shot metric-based methods show better interpretability than fine-tuned ones.
@@
However, DetectGPT still has three significant defects: (i) DetectGPT’s reliance on the logit regression module’s threshold compromises its generalization in zero-shot settings and limited to large batch input, failing on individual inputs. (ii) DetectGPT does not fully utilize the perturbation. As a metrics-based method, it only considers the probability difference caused by perturbation, which is overly simplified and slightly indistinguishable. Perturbation should indeed be a stronger augment that carries implicit language pattern information. (iii) DetectGPT perturbs the original sample randomly and unrestricted, which could introduce more noise and negatively impact the performance Kim et al. (2022). For example, Liu et al. (2023) find entity-relationship plays a role in the detection, which might be destroyed in random perturbation of DetectGPT.
In this paper, we thus propose a Perturbation-based Contrastive Learning model, Pecola, for MGT detection, toward the defects via two stages, i.e., Selective Strategy Perturbation (Sec. 3.1) and Token-Level Weighted Multi-Pairwise Contrastive Learning (Sec. 3.2). Firstly, Selective Strategy Perturbation is a token-level rewriting method with restrictions on modifying important texts Campos et al. (2020) to reduce noise. The motivation is to simulate the human behavior of modification Verma and Lee (2017); Fetaya et al. (2020); Wang et al. (2019). The perturbation strategy consists of token removal and substitution, as shown in Fig. 1. The experiments show that the Selective Strategy Perturbation method can improve the performance of both metrics-based (i.e., DetectGPT) and model-based methods. Secondly, we propose a Multi-Pairwise Contrastive Learning model to process the perturbed texts. Different from the logit regression module in DetectGPT, the trained model is generalizable without any threshold setting, and it can deal with individual inputs. Moreover, by utilizing multi-pairwise contrastive learning, the model could better utilize perturbation to focus on the language pattern gap between HWT and MGT. The importance weight from the perturbation stage is also reused as contrastive learning weight. Notably, by using contrastive learning, Pecola is a strong few-shot fine-tuning method, which effectively bridges and integrates metric-based and fine-tuned detector categories. Finally, extensive experiments show Pecola is significantly superior to baseline and SOTA methods on four datasets, Pecola improves by 1.20% to SOTA on average under few-shot settings, surpassing the latest methods by 3.84% among metric-based detectors and by 1.62% among fine-tuned detectors. Further experiments show that Pecola is also better at generalization, robustness, and effectiveness.
@@
Outperformance: Our detector Pecola outperforms all eight compared models on four public datasets. And Pecola is more robust to the choice of base model and filling model. Furthermore, we prove its generalization ability across domains and generators of data.
@@ 2 Related Work
Machine-generated Text Detection. While fine-tuned detectors have proven effective for MGT detection Wahle et al. (2022); Hu et al. (2023), the requirement for annotated datasets poses a significant challenge due to the proliferation of unchecked, high-quality generated texts. To address this challenge, DetectGPT Mitchell et al. (2023) and Fast-DetectGPT Bao et al. (2024) have demonstrated strong performance in white-box zero-shot settings. Similarly, CoCo Liu et al. (2023) is designed to detect MGT with low resource annotations, utilizing a coherence-based contrastive learning model. Moreover, SeqXGPT Wang et al. (2023) utilize log probability lists from white-box LLMs as features; Sniffer Shi et al. (2024) and GPT-Who Venkatraman et al. (2023) place more emphasis on tracing the origin of MGT. Recently, watermarking Kirchenbauer et al. (2023) is introduced to mitigate the risk associated with unchecked MGTs by embedding imperceptible signals within text outputs during generation. In contrast to previous methods, our approach integrates data perturbation with contrastive learning, placing particular emphasis on reducing reliance on mask-filling models and enhancing performance in few-shot scenarios.
@@
Perturbation. Data perturbation methods find frequent application in text classification tasks Gao et al. (2022); Shum et al. (2023), which is commonly employed through the technique of consistency regularization Xie et al. (2020); Chen et al. (2020). Nevertheless, in MGT detection, previous perturbation methods have exhibited certain limitations. For instance, they often resort to randomly selecting target tokens for synonym replacement Wang et al. (2018), deletion, insertion Wei and Zou (2019), rewriting by LLMs Mao et al. (2024), and fine-tuning pre-trained language models (PLMs) to fill text spans of variable lengths Gao et al. (2022). While these methods do enhance text diversity, the indiscriminate replacement of tokens without guided rules can lead to the generation of less reliable texts. Wang et al. (2024) utilize perturbations as stress test approaches for the robustness of MGT detectors to show their loopholes. These limitations motivate us to devise data perturbation methods tailored for MGT detection. Our approach, with selective perturbation, aims to better represent meaningful recombination spaces while preserving the inherent semantic features of the text, ultimately enhancing the diversity of samples.
@@
Constrastive Learning. Contrastive learning is an effective solution method to the issues that solely relying on cross-entropy classification loss would lead to a lack of robustness and suboptimal generalization Tack et al. (2020); Hu et al. (2023). In limited labeled data task Gunel et al. (2021), introduce a robust contrastive learning method to capture the similarities between the same instances in the representation space while separating those of different classes. Similarly, out-of-distribution (OOD) usually leads to severe semantic shift issues during inference, prompting another approach based on margin contrastive learning Zhou et al. (2021). Differently, our method focuses more on the changes of the rephrase space in data distribution after perturbation, and strives to reduce reliance on the mask-filling models in few-shot learning.
@@ 3 Methodology
As shown in Fig. 2, the workflow of Pecola mainly consists of two stages: Selective Strategy Perturbation and Supervised Contrastive Learning, which joined the advantage of metric-based and model-based detection methods, respectively.
@@ 3.1 Selective Strategy Perturbation
Token Importance Assessment. To accurately assess the significance of tokens within the text and mitigate information loss stemming from random masking, we expand upon the YAKE algorithm Campos et al. (2020) to operate at the token level. The YAKE algorithm builds upon certain assumptions Machado et al. (2009), which posit that the importance of a candidate word decreases as the richness of the vocabulary surrounding it increases. This fundamental assumption remains applicable when processing text at the token level, i.e., token importance assessment.
@@
Mask Position Selection. After getting the important tokens set of each text input , we use special token [MASK] to replace some of the tokens in the text input to construct masked perturbation input . In order to relieve the information loss caused by masking perturbation, we add regularization to the vanilla random masking method and use a selective masking strategy to prevent important tokens from being masked.
@@
where is the mask ratio, and represents an indicator function with a value of if and only if the condition is satisfied, otherwise, it is . Then we gather all masked perturbation inputs and include them in the training set to give the model masked perturbation, improving model robustness.
Mask-Filling. Additionally, we utilize PLMs, e.g., T5 Raffel et al. (2020) or RoBERTa Liu et al. (2019) etc., to fill the masked perturbation inputs and create the filled perturbation inputs . Similar to the above, we include all filled perturbation inputs in the training set and obtain the final training set .
@@ 3.2 Token-Level Weighted Multi-Pairwise Contrastive Learning
This adaptive margin ensures that the model is steered to maintain discriminative embeddings despite data perturbation during training. Then we get the following contrastive loss as:
@@
| Dataset | Metric | Shot | RoBERTa | GLTR† | CE+SCL | CE+Margin | IT:Clust | CoCo | DetectGPT | Fast-Detect. | Pecola |
@@ 4.2 Comparison Models
RoBERTa Liu et al. (2019), supervised methods via standard fine-tuning PLMs as classifiers. We use RoBERTa-base (125M).
GLTR Gehrmann et al. (2019), a metric-based detector and based on next-token probability. We follow the setting of Guo et al. (2023), utilizing the Test-2 feature. For a fair comparison with fine-tuning methods, we first use the few-shot training samples to settle the threshold and adapt the fixed threshold in the test set.222The base model of GLTR is chosen based on the generator of the dataset: for GPT-2 and Grover datasets, we use GPT-2 Small (124M); and for GPT-3.5 and HC3 datasets, we use GPT-J (6B) Wang (2021), which is the best open-source model to simulate ChatGPT and GPT-3.5 empirically.
@@
IT:Clust Shnarch et al. (2022), a general text classification method that employs unsupervised clustering as an intermediate for fine-tuning PLMs, utilizing RoBERTa-base.
@@
DetectGPT Mitchell et al. (2023), a zero-shot metric-based MGT detector, using T5-large Raffel et al. (2020) to perturb texts. Same as GLTR, we fix the threshold.333For all four datasets (including HC3 and GPT-3.5 datasets), we use GPT-2 Small (124M) as the base model to calculate the likelihood. The reason is Mireshghallah et al. (2023) find that small model is better black-box detector for DetectGPT.
@@ 4.3 Performance Comparison
Moreover, the unsupervised learning methods tend to show better performance in extremely few shot scenarios. Unsurprisingly, unsupervised methods do not see a notable performance improvement with the increase in the number of training samples, which causes them to outperform on the fewest shot settings initially but soon be surpassed. As for the deception of generators, Grover appears to be the hardest to detect, while other models are relatively “honest” to detectors. It might have originated from the adversarial training strategy of Grover, while the bulit-in detector module adversarially shifts the LLM’s detectable features. More interestingly, advanced language models show a weaker ability to cheat detectors. Most detectors achieve around 98% in accuracy on the GPT-3.5 and HC3 datasets, which is consistent with the conclusion from Liu et al. (2023); Chen et al. (2023). We hypothesize that the easy-to-detect nature may originate from the lack of semantics diversity in GPT-3.5 and ChatGPT as they use RLHF Kirk et al. (2023).
@@ 4.4 Ablation Study
To illustrate the effectiveness of the Pecola components, we do the ablation experiments on the Selective Strategy Perturbation stage and the Contrastive Learning stage on the 64-example GPT-2 dataset. We also demonstrate the Scalability of Pecola in Appendix C.1.
| Method | Acc | F1 | |
| w/o. mask | 78.00 | 77.93 | |
| w/o. mask-fill | 77.78 | 77.72 | |
| w/o. mask. CLw | 75.80 | 75.23 | |
| w/o. mask-fill. CLw | 75.56 | 75.10 | |
| w/o. CLw | 76.60 | 76.22 | |
| w/o. | 78.02 | 77.93 | |
@@
Ablation on Contrastive Learning. It primarily investigates the impact of CE and contrastive loss. (i) w/o. CLw refers to the model ablating weighted contrastive learning; (ii) w/o. refers to the model including contrastive learning but ablating weight.
@@ 4.5.1 Model Qualities
Analysis on Robustness. To validate the robustness of Pecola in the few-shot learning settings, we apply four post hoc perturbation operations for each token in the test dataset randomly, i.e., deletion, replacement, insertion, and repetition. As indicated in Table 3, for each perturbation method employed, our decline rate is consistently lower compared to the baseline RoBERTa. On average, Pecola maintains a 5.66% higher accuracy and an 8.77% superior F1-score. Specifically, in the deletion method, where we introduce a 15% random perturbation, it is noteworthy that the accuracy of Pecola decreases merely 1.64%, underscoring its remarkable robustness.
| Model | RoBERTa | Pecola | |||
| Metric | Acc | F1 | Acc | F1 | |
| Original | 74.412.47 | 73.912.69 | 78.921.14 | 78.881.17 | |
| Delete | 71.775.88(-2.640) | 70.428.05(-3.490) | 77.281.79(-1.640) | 77.062.03(-1.820) | |
| Repeat | 64.696.63(-9.720) | 61.749.20(-12.17) | 69.744.83(-9.180) | 67.876.24(-11.01) | |
| Insert | 50.750.67(-23.66) | 36.441.60(-37.47) | 57.612.52(-21.31) | 49.294.57(-29.59) | |
| Replace | 52.041.58(-22.37) | 39.483.59(-34.43) | 57.252.21(-21.67) | 48.893.93(-29.99) | |
| Average | 59.81 (-14.60) | 52.02 (-21.89) | 65.47 (-13.45) | 60.78 (-18.10) | |
Analysis on Affinity. Affinity pertains to alterations in data distribution resulting from perturbations, quantified by observing the fluctuations in accuracy. We demonstrate the superiority of the selective masking method over the random masking method using the Affinity metric, following the setting of DetectGPT. We applied a 15% mask proportion with a span of 2 tokens on the test dataset and simultaneously employed T5-Large Raffel et al. (2020) as the mask-filling model. We trained RoBERTa-base and Pecola on the 64-example GPT2 dataset. As shown in Table 4, in comparison to the random masking perturbation method utilized in DetectGPT, we observe a 1.92% and 0.49% increase in Affinity when employing the selective masking method. Additionally, the mask-filling method yields affinity improvements of 3.38% and 1.32% for RoBERTa and Pecola models, respectively. These results illustrate that the Selective Multi-Strategy Perturbation method effectively preserves more distinguishable features between MGTs and HWTs.
| Model | RoBERTa | Pecola | |
| Random Mask | -2.64 | -1.64 | |
| Selective Mask | -0.72 | -1.15 | |
| Mask-Filling | -4.72 | -2.66 | |
@@
Analysis on Diversity Conversely, diversity assesses the range and variability of perturbed data, utilizing metrics Dist-1 and Dist-2 Celikyilmaz et al. (2020). Here, we use three common perturbation methods to demonstrate the importance of not arbitrarily changing important tokens and the significance of select masks. (1) Token Substitution (TS, Zhang et al. 2015), replaces tokens with synonyms from WordNet Miller (1992); (2) SwitchOut (SO, Wang et al. 2018), uniformly samples and randomly substitutes from the vocabulary of test samples; and (3) Two-stage (TWs, Wei et al. 2021) trains the mask-filling model on the original data.
The ideal perturbation result is to have high Affinity scores while ensuring high Diversity scores Celikyilmaz et al. (2020). As shown in Table 5, through Selective Strategy Perturbation, models achieve better diversity with high distribution shifts. And the overall improvement in Affinity by over 18% also shows greater diversity than the original data. The above results demonstrate the superiority of our perturbation method.
| Method | Affinity | Dist-1 | Dist-2 | |||
| TS | -20.00 | 3.38 | 43.43 | |||
| TO | -22.06 | 6.81 | 53.61 | |||
| TWs | -21.13 | 3.24 | 41.85 | |||
| Original | - | 8.70 | 50.32 | |||
| Pecola | -1.34 | 15.59 | 57.01 |